# In what ways do text to image models fail?

### v5

@@ 3.1 Human and Animal Body Parts
Limbs, Hands, and Fingers. The models used for generating deepfakes often fall short when it comes to accurately depicting the intricate details of human extremities. For instance, hands may randomly duplicate, fingers can merge together or there may be too many or too few of them, and third legs may unexpectedly appear while existing limbs may disappear without a trace. Furthermore, limbs may be positioned in unrealistic or impossible poses, or there may be an excess number of them. As a result, deepfakes may exhibit unnatural body language, such as unrealistic gestures or postures that are out of place. See Figs 11 and 12.
@@ 3.4 Semantics and Logic
Spatial Reasoning. Natural scenes are complex and contain a wide range of spatial relationships among objects, such as occlusions, relative distances, and orientations. Capturing these relationships requires the model to have a nuanced understanding of the scene and the objects within it, which can be difficult to achieve without more explicit guidance. Furthermore, some image generation models rely solely on pixel-level reconstruction, without explicitly modeling the underlying semantics or spatial relationships. In these cases, the model may generate images that are visually realistic but lack coherent semantic meaning or accurate spatial relationships among objects. Please see Figure 23 for some examples.
@@ 3.5 Text, Noise, and Details
Accurately rendering all details in complex scenes or crowd scenes, such as those depicted in Figures 31 and 32, can be particularly challenging for AI. The complexity of these scenes makes it difficult for the AI to accurately model every detail and can lead to errors in object placement, lighting, perspective, and other features. Despite the challenges, AI technology continues to improve, and advancements are being made in the generation of more realistic and believable large and crowd scenes.
@@ 4.2 Some Challenging Objects
Generative models face particular challenges when it comes to generating images of objects such as clocks, Lego houses, chessboards, carpets, circuit boards, basketballs, glasses of water, dice, diagrams and tables, keyboards, and computer screens. One of the reasons for this is that these types of images contain many repeated patterns, which can be difficult for the model to accurately capture. Several examples of failed attempts to generate these objects can be seen in Figures 33 and 34. This list of challenging objects can be used to assess and compare the performance of different image generation models.

### v6

# Qualitative Failures of Image Generation Models and Their Application in Detecting Deepfakes

@@ 3.1 Human and Animal Body Parts
Limbs, Hands, and Fingers. The models used for generating deepfakes often fall short when it comes to accurately depicting the intricate details of human extremities. For instance, hands may randomly duplicate, fingers can merge together or there may be too many or too few of them, and third legs may unexpectedly appear while existing limbs may disappear without a trace. Furthermore, limbs may be positioned in unrealistic or impossible poses, or there may be an excess number of them. As a result, deepfakes may exhibit unnatural body language, such as unrealistic gestures or postures that are out of place. In certain instances, models are unable to accurately depict the interaction between objects and body parts (e.g. brushing). See Figs 11 and 12.
@@ 3.4 Semantics and Logic
Spatial Reasoning. Natural scenes are complex and contain a wide range of spatial relationships among objects, such as occlusions, relative distances, and orientations. Capturing these relationships requires the model to have a nuanced understanding of the scene and the objects within it, which can be difficult to achieve without more explicit guidance. Furthermore, some image generation models rely solely on pixel-level reconstruction, without explicitly modeling the underlying semantics or spatial relationships. In these cases, the model may generate images that are visually realistic but lack coherent semantic meaning or accurate spatial relationships among objects. Please see Figures LABEL:fig:reasoning,fig:promptfailure for some examples.
@@
Figure 24: Examples for which generative models do not understand the propmpt properly. Left image is generated using ChatGPT which is using DALL-E 3 in the background. Similar phenomenon was observed using Bing image creator. Right image is generated by Google Gemini.
@@ 3.5 Text, Noise, and Details
Accurately rendering all details in complex scenes or crowd scenes, such as those depicted in Figures 32 and 33, can be particularly challenging for AI. The complexity of these scenes makes it difficult for the AI to accurately model every detail and can lead to errors in object placement, lighting, perspective, and other features. Despite the challenges, AI technology continues to improve, and advancements are being made in the generation of more realistic and believable large and crowd scenes.
@@ 4.2 Some Challenging Objects
Generative models face particular challenges when it comes to generating images of objects such as clocks, Lego houses, chessboards, carpets, circuit boards [^5] In situations like this, zooming in on the image will highlight the deficiencies and blurred regions., basketballs, glasses of water, dice, diagrams and tables, keyboards, and computer screens. One of the reasons for this is that these types of images contain many repeated patterns, which can be difficult for the model to accurately capture. Several examples of failed attempts to generate these objects can be seen in Figures 34 and 35. This list of challenging objects can be used to assess and compare the performance of different image generation models.
@@ 4.4 Failure modes from other studies
Generative image models have also problems with bias and discrimination, similar to LLMs [7]. People discovered that requesting Google Gemini to generate images of certain historical events or figures led to amusing outcomes. For example, the Founding Fathers, historically known as white slave owners, were depicted as a multicultural group that included people of color (see https://techcrunch.com/2024/02/23/embarrassing-and-wrong-google-admits-it-lost-control-of-image-generating-ai/). Please see Figure 39.
Figure 39: An example of bias failure, generated by Google Gemini.
For additional failures reported by the general public using generative image models, please see https://www.boredpanda.com/ai-fails/.